# Human strategic decision making in parametrized games

### 23 Mar 2026 v4

@@ 2 Parametrized games
A strategy-parametrized strategic-form game is a tuple $(\{G_{\mathbf{\lambda}}\},\Lambda)$ where for each real-valued vector of parameters $\mathbf{\lambda}\in\Lambda$, $G_{\mathbf{\lambda}}$ is a tuple $(N,S_{\mathbf{\lambda}},u)$ with finite set of players NN, finite set of pure strategies $S_{\mathbf{\lambda},i}$ for each player $i\in N$, and real-valued utility for each player for each strategy vector (aka strategy profile), $u_{i}:\times_{j}S_{\mathbf{\lambda},j}\rightarrow\mathbb{R}$.
@@
While the strategic form can be used to model simultaneous actions, another representation, called the extensive form, is generally preferred when modeling settings that have sequential moves. The extensive form can also model simultaneous actions, as well as chance events and imperfect information (i.e., situations where some information is available to only some of the agents and not to others). Extensive-form games consist primarily of a game tree; each non-terminal node has an associated player (possibly chance) that makes the decision at that node, and each terminal node has associated utilities for the players. Additionally, game states are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. A pure strategy for player $i$ is a mapping that selects an action at each information set belonging to player ii.
@@ 6.2 Generalized Kuhn Poker
  • Player AA acts first and can either bet \$1 or check
    • If AA bets, player $B$ can call or fold
        If AA bets and $B$ calls, then whoever has the higher card wins the \$4 pot
        If AA bets and $B$ folds, then AA wins the entire \$3 pot
    • If AA checks, $B$ can bet \$1 or check.
        If AA checks and $B$ bets, then AA can call or fold.
          If AA checks, $B$ bets, and AA calls, then whoever has the higher card wins the \$4 pot
          If AA checks, $B$ bets, and AA folds, B wins the \$3 pot
        If AA checks and $B$ checks, then whoever has the higher card wins the \$2 pot
An analysis of the equilibria is provided in the appendix. The equilibrium strategies contain some elements of deceptive behavior, as is present in larger variants of poker. For example, player AA sometimes checks with a 3 as a trap or slowplay, and both players sometimes bet with a 1 as a bluff.
@@ Appendix C Generalized Kuhn Poker
  • Player AA acts first and can either bet \$1 or check
    • If AA bets, player $B$ can call or fold
        If AA bets and $B$ calls, then whoever has the higher card wins the \$4 pot
        If AA bets and $B$ folds, then AA wins the entire \$3 pot
    • If AA checks, $B$ can bet \$1 or check.
        If AA checks and $B$ bets, then AA can call or fold.
          If AA checks, $B$ bets, and AA calls, then whoever has the higher card wins the \$4 pot
          If AA checks, $B$ bets, and AA folds, then B wins the \$3 pot
        If AA checks and $B$ checks, then whoever has the higher card wins the \$2 pot
@@
  • If AA bets in the first round, then:
@@
  • If AA checks in the first round, then:
@@
  • If AA checks and $B$ bets, then:
@@
    Equilibrium strategies contain some elements of deceptive behavior. For example, player AA sometimes checks with a 3 as a trap or slowplay, and both players sometimes bet with a 1 as a bluff.

### 24 Mar 2026 v5

@@ 2 Parametrized games
A strategy-parametrized strategic-form game is a tuple $(\{G_{\mathbf{\lambda}}\},\Lambda)$ where for each real-valued vector of parameters $\mathbf{\lambda}\in\Lambda$, $G_{\mathbf{\lambda}}$ is a tuple $(N,S_{\mathbf{\lambda}},u)$ with finite set of players $N$, finite set of pure strategies $S_{\mathbf{\lambda},i}$ for each player $i\in N$, and real-valued utility for each player for each strategy vector (aka strategy profile), $u_{i}:\times_{j}S_{\mathbf{\lambda},j}\rightarrow\mathbb{R}$.
@@
While the strategic form can be used to model simultaneous actions, another representation, called the extensive form, is generally preferred when modeling settings that have sequential moves. The extensive form can also model simultaneous actions, as well as chance events and imperfect information (i.e., situations where some information is available to only some of the agents and not to others). Extensive-form games consist primarily of a game tree; each non-terminal node has an associated player (possibly chance) that makes the decision at that node, and each terminal node has associated utilities for the players. Additionally, game states are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. A pure strategy for player $i$ is a mapping that selects an action at each information set belonging to player $i$.
@@ 6.2 Generalized Kuhn Poker
  • Player $A$ acts first and can either bet \$1 or check
    • If $A$ bets, player $B$ can call or fold
        If $A$ bets and $B$ calls, then whoever has the higher card wins the \$4 pot
        If $A$ bets and $B$ folds, then $A$ wins the entire \$3 pot
    • If $A$ checks, $B$ can bet \$1 or check.
        If $A$ checks and $B$ bets, then $A$ can call or fold.
          If $A$ checks, $B$ bets, and $A$ calls, then whoever has the higher card wins the \$4 pot
          If $A$ checks, $B$ bets, and $A$ folds, B wins the \$3 pot
        If $A$ checks and $B$ checks, then whoever has the higher card wins the \$2 pot
An analysis of the equilibria is provided in the appendix. The equilibrium strategies contain some elements of deceptive behavior, as is present in larger variants of poker. For example, player $A$ sometimes checks with a 3 as a trap or slowplay, and both players sometimes bet with a 1 as a bluff.
@@ Appendix C Generalized Kuhn Poker
  • Player $A$ acts first and can either bet \$1 or check
    • If $A$ bets, player $B$ can call or fold
        If $A$ bets and $B$ calls, then whoever has the higher card wins the \$4 pot
        If $A$ bets and $B$ folds, then $A$ wins the entire \$3 pot
    • If $A$ checks, $B$ can bet \$1 or check.
        If $A$ checks and $B$ bets, then $A$ can call or fold.
          If $A$ checks, $B$ bets, and $A$ calls, then whoever has the higher card wins the \$4 pot
          If $A$ checks, $B$ bets, and $A$ folds, then B wins the \$3 pot
        If $A$ checks and $B$ checks, then whoever has the higher card wins the \$2 pot
@@
  • If $A$ bets in the first round, then:
@@
  • If $A$ checks in the first round, then:
@@
  • If $A$ checks and $B$ bets, then:
@@
    Equilibrium strategies contain some elements of deceptive behavior. For example, player $A$ sometimes checks with a 3 as a trap or slowplay, and both players sometimes bet with a 1 as a bluff.